OpenStudy (anonymous):

What is the solution of the equation 9 = -2x + 15?

OpenStudy (texaschic101):

9 = -2x + 15 --- subtract 15 from both sides 9 - 15 = -2x -6 = -2x -- divide by -2 -6/-2 = x 3 = x

OpenStudy (texaschic101):

any questions ?